A sample of an ideal gas is isothermally expanded at 47.0 ℃ and the gas does 575 J of work during this expansion. Calculate the change in the entropy of the gas.
Answer #1

From first law of thermodynamics

Q= delta U+ W

in isothermal process delta U=0

dQ= dW= 575 J

T= 273+47= 320 K

dS= dW/T

dS= 575/320= 1.7968 J/K

change in entropy is 1.7968 J/K

Thank you